# DemoDiffusion: One-Shot Human Imitation using pre-trained Diffusion Policy

### v1

@@ 4 Experiments
4.1 Dexterous Grasping in Simulation
@@
For training the robot policy, we collect total 985 grasping trajectories of Allegro hand over 58 training objects[50] (randomly sampled 26 and 32 objects from ShapeNet [51] and PartNet [52], respectively), resulting in 22.2 success rate. We use a variant of 3D Diffusion Policy [53], which takes 3D point tracks of an object instead of point clouds. At test time, we provide human grasping trajectories on a subset of Objvaerse dataset [54], which are unseen during training. The test set contains total 1220 objects, with 1 human grasping trajectory per object, provided by GraspXL [50]. Note that there is no error from hand estimation or the initial configuration of the object in this setup. We use ground-truth human hand 3D keypoints, and the object is located at the same location as the human demonstration.
Figure 4: Real-World Manipulation Tasks. Each column corresponds to the respective human demonstrations for the real-world manipulation tasks, showing two frames from each demonstration.
Figure 5: Qualitative comparisons for real-robot manipulation. We show rollouts from two tasks for DemoDiffusion and the baselines. The respective strips show progression of rollouts from left to right including the start frame, an intermediate frame, and the final frame. Frames are cropped for better visibility. Please refer to the website for detailed visualizations.
Results.
As in GraspXL [50], we group the objects based on their size, resulting in small, medium, and large objects. We compare DemoDiffusion with two baselines. The first is open-loop kinematic retargeting obtained from the human demonstration. The second is directly deploying the robot policy we trained.
Table. 1 shows the main results. Our method outperforms baselines, with more performance gain for small-sized objects. We hypothesize this roots from that larger objects are easier to grasp, supported by robot policy performance improving for larger sized objects. This experiment shows that even in a highly controlled setup where the robot environment is initialized identically to that of human demonstration, kinematic retargeting cannot solve the task, while slightly refining with a generalist grasping policy significantly improves the performance. Using kinematic retargeting to initiate the denoising process also helps the denoising process itself, outperforming the robot policy. Additionally, DemoDiffusion achieves higher inference speed, by the factor of $S/s^{*}$.
Ablation studies.
We analyze the influence of $s^{*}$ on DemoDiffusion in Fig. 3. Overall, we observe a consistent trend of performance gain from $s^{*}/S=1$ to $s^{*}/S=0.2$, showing that DemoDiffusion is robust to the choice of hyperparameter.
4.2 Real-World Manipulation
@@
Figure 6: Workspace with 5 cameras. We use the four external cameras for triangulation to obtain the global pose of the hand mesh from a human demonstration. The pre-trained diffusion policy uses the two cameras marked in purple.
@@
For the real-world experiments, we use a Franka Emika Panda arm equipped with a two-finger gripper from Robotiq. We use the pre-trained diffusion policy $\bar{\pi}_{\theta}(\mathbf{a}_{t}|\mathbf{o}_{\leq t},\mathcal{T})$ called Pi-0 released by Physical Intelligence [1]. This policy takes as input a language task specification $\mathcal{T}$, and observations from two cameras in the scene, and outputs per-timestep joint velocity. This policy was trained on a large offline dataset of tele-operated robot demonstrations, and is a generalist policy that can perform a broad set of tasks. We do not fine-tune or adapt this policy in any way and directly use it for DemoDiffusion. We use an off-the-shelf monocular hand reconstruction model, Hamer [47] to obtain per-timestep hand mesh reconstructions from each of the external cameras given a human demonstration in the scene. Our experiments follow the protocol of a human first demonstrating an object manipulation in the scene, followed by robot execution after re-setting the scene.
Workspace.
Fig. 6 describes our real-robot manipulation setup. The scene has 5 cameras including 4 Real-sense external cameras and 1 Zed-mini wrist camera mounted on the robot gripper. We calibrate the four external cameras and use them for triangulating the 3D hand pose so that we can obtain the position and orientation of a human hand in the world coordinates, with origin at the base of the robot. The pre-trained diffusion policy Pi-0 requires just two cameras (marked in purple) and the policy does not use any calibration information.
Tasks.
We perform evaluations for 8 different manipulation tasks, the language description and the human demonstration for which are described in Fig. 4. Each of these tasks require a different manipulation strategy involving either prehensile or non-prehensile manipulation. We collect just a single human demonstration per task but for robot execution, we make organic variations in the location of objects in the scene. Additionally, we add other objects in the scene as visual distractors, in order to see whether DemoDiffusion can generalize under such settings. All variations are consistent across the baselines and our method. Please refer to the Appendix for more explanations.
Baselines.
For the real-world experiments we compare DemoDiffusion with two baselines. The first is open-loop kinematic retargeting obtained from the human demonstration. The second is directly deploying the Pi-0 model with the language instruction for the respective task. Since Pi-0 operates in the join velocity space, for DemoDiffusion we first convert the kinematically retargeted trajectory to the joint velocity space with IK and then add noise to the trajectory, followed by denoising with Pi-0. We deploy DemoDiffusion with a heuristically selected noise level $s^{*}/S$ depending on the task – preferring the kinematic trajectory with $s^{*}/S=0.2$ when Pi-0 has a zero success rate and $s^{*}/S=0.4$ otherwise. Note that in the limit of 00 noise added, we recover the first baseline, kinematic retargeting, and in the limit of maximum noise $1$ we obtain the second baseline Pi-0.
Qualitative Visualizations.
In Fig. 5 we show rollouts for two tasks across all the methods. We observe that Pi-0 is only able to perform general reaching behaviors but doesn’t follow the required manipulation trajectory necessary to perform these tasks, or confuse the target object to manipulate. On the other hand, we observe that (open-loop) kinematic retargeting from the human demonstration has a plausible motion of the robot end-effector, but it completely misses grasping the iron, and loses contact with the laptop before fully closing it. Compared to these, DemoDiffusion is able to reach the respective object and also follow the approximate motion of the retargeted trajectory while remaining in contact with the object in order to solve the tasks.
Quantitative Comparisons.
In Table 2 we compare DemoDiffusion and the baselines across all the real-world manipulation tasks. We observe that DemoDiffusion either performs comparably or outperforms all the baselines in all the tasks. Interestingly, even for tasks like closing the laptop where both the baselines fail, DemoDiffusion is able to leverage the best of both the baselines (that represent extremes of our approach under the hyper-parameter $s^{*}$) in order to achieve a non-zero success rate.
@@
Table 2: Quantitative Results for Real-World Robot Manipulation. Success rates (%) over 10 trials per task. Please refer to the Appendix for more comprehensive quantitative comparisons.
Method
Shut
Laptop
Close
Microwave
Drag
Basket
Wipe
Table
Iron
Curtain
Pick up
Bear
Pick Place
Banana
Pick Place
Bowl
Avg.
Pi-0 20 10 10 50 20 0 0 0 13.8
Kinematic Retargeting 10 80 80 0 70 40 70 70 52.5
DemoDiffusion 60 90 80 100 90 60 90 100 83.8
5 Discussion
@@
We presented DemoDiffusion, a simple approach to leverage generalist diffusion models for imitating a human demonstration. By retargeting hand motion from the human demonstration into an approximate robot end-effector trajectory and modifying it through de-noising with a pre-trained diffusion policy, DemoDiffusion enables the robot to produce actions that are both aligned with the demonstrated intent and contextually plausible. Our method consistently outperforms both the base diffusion policy and the raw retargeted trajectory across diverse tasks in simulation and the real world, without requiring additional training, paired demonstrations, or reward functions. We envision that this simple approach can serve as a starting point for future efforts at human imitation across generic scenarios and perhaps also yield an improved exploration strategy for methods pursuing policy adaptation with online RL.
@@
Our approach has several limitations. First, it assumes that the robot should act similarly to the human in order to successfully complete the task, which may not hold in scenarios requiring different strategies due to embodiment or environmental constraints. Second, while effective for one-shot imitation (and not requiring and training / test-time fine-tuning), the method does not produce a reusable task policy that can generalize across arbitrary variations of the task. Third, the quality of the retargeted trajectory is crucial: accurate 3D human motion capture is challenging and errors in retargeting can impact downstream performance. Additionally, the method implicitly assumes that the timing and speed of human and robot actions are aligned; extending the approach to allow for temporal alignment at test time is a promising direction of future work.
@@ Acknowledgements
We appreciate the helpful discussions with Yanbo Xu, Qitao Zhao, and Lucas Wu. We would also like to thank Kenny Shaw, Tony Tao, Jiahui Yang, Andrew Wang, Jason Liu, Hengkai Pan, Mohan Kumar Srirama for helping setting up the hardware. This work was supported by gift awards from CISCO and Google.
@@ A.1 Hardware Setup
A.2 Task Setup Description
@@
A.4 Inference Details
At each timestep $t$, we index the human demonstration at same timestep $t$, and compute the kinematically retargeted actions. As Pi-0 predicts a chunk of actions with length $H=10$, we also use the action chunk computed from human demonstration from timestep $t$ to $t+H-1$. The episode length is set to be the same as the length(number of timesteps) of human demonstrations across all methods.
A.5 Kinematic Retargeting Implementation Details
@@ A.6 Hyperparameters
Table 3: Hyperparameters in Real-world.
@@
A.7 Additional Results
Results with Different Hyperparameters. We present the results of DemoDiffusion using two different hyperparameter settings, $s^{*}/S=0.2$ and $s^{*}/S=0.4$. The results are summarized in Table 4. Overall, we observe a consistent trend based on the performance of the base diffusion policy (Pi-0):
@@ A.7 Additional Results
Table 4: Additional Quantitative Results for Real-World Robot Manipulation.
Method
Shut
Laptop
Close
Microwave
Drag
Basket
Wipe
Table
Iron
Curtain
Pick up
Bear
Pick Place
Banana
Pick Place
Bowl
Avg.
Pi-0 20 10 10 50 20 0 0 0 13.8
Kinematic Retargeting 10 80 80 0 70 40 70 70 52.5
DemoDiffusionw$(s^{*}/S=0.2)$ 40 80 70 60 70 60 90 100 71.3
DemoDiffusionw$(s^{*}/S=0.4)$ 60 90 80 100 90 40 20 0 60.0
@@ Appendix B Dexterous Grasping in Simulation
B.1 Task Setup Description
@@
B.2 Kinematic Retargeting Implementation Details
@@
B.3 Robot Policy Implementation Details
@@ B.4 Hyperparameters
Table 5: Hyperparameters in Simulation.

### v2

@@ IV Experiments
  • How robust is DemoDiffusion to noise in human demonstration, or different kinematic retargeting?
IV-A Dexterous Grasping in Simulation
@@
For training the robot policy, we collect a total of 985 grasping trajectories of Allegro hand over 58 training objects[59] (randomly sampled 26 and 32 objects from ShapeNet [7] and PartNet [38], respectively), resulting in 22.2% success rate. We use a variant of 3D Diffusion Policy [58], which takes 3D point tracks of an object instead of point clouds. At test time, we provide human grasping trajectories on a subset of Objaverse dataset [13], which are unseen during training. The test set contains 1220 objects, with 1 human grasping trajectory per object, provided by GraspXL [59]. Note that there is no error from hand estimation or the initial configuration of the object in this setup. We use ground-truth human hand 3D keypoints, and the object is located at the same location as the human demonstration.
@@
Figure 5: Workspace with 5 Cameras. We use the four external cameras for triangulation to obtain the global pose of the hand mesh from a human demonstration. The pre-trained policy uses the two cameras marked in purple.
Results. As in GraspXL [59], we group the objects based on their size, resulting in small, medium, and large objects. We compare DemoDiffusion with two baselines. The first is open-loop kinematic retargeting obtained from the human demonstration. The second is directly deploying the robot policy we trained.
Table I shows the main results. Our method outperforms baselines, with the largest gains on small objects. We hypothesize this is because larger objects are easier to grasp, supported by base robot policy performance improving for larger objects. This experiment shows that even in a highly controlled setup where the robot environment is initialized identically to that of human demonstration, kinematic retargeting cannot solve the task, while slightly refining with a generalist grasping policy significantly improves the performance. Using kinematic retargeting to initiate the denoising process also helps the denoising process itself, outperforming the robot policy. Additionally, DemoDiffusion achieves higher inference speed, by the factor of $S/s^{*}$.
Ablation studies. We analyze the influence of $s^{*}$ on DemoDiffusion in Fig. 3. Overall, we observe a consistent trend of performance gain from $s^{*}/S=1$ to $s^{*}/S=0.2$, showing that DemoDiffusion is robust to the choice of hyperparameter.
IV-B Real-World Manipulation
@@
For the real-world experiments, we use a Franka Emika Panda arm equipped with a two-finger gripper from Robotiq. We use the publicly available pre-trained flow-matching policy $\bar{\pi}_{\theta}(\mathbf{a}_{t}|\mathbf{o}_{\leq t},\mathcal{T})$ called Pi-0 released by Physical Intelligence [5], trained on the DROID dataset [26]. This policy takes as input a language task specification $\mathcal{T}$, and observations from two cameras in the scene, and outputs per-timestep joint velocity. This policy was trained on a large offline dataset of tele-operated robot demonstrations, and is a generalist policy that can perform a broad set of tasks. We do not fine-tune or adapt this policy in any way and directly use it for DemoDiffusion. We use an off-the-shelf monocular hand reconstruction model, Hamer [43] to obtain per-timestep hand mesh reconstructions from each of the external cameras given a human demonstration in the scene. Our experiments follow the protocol of a human first demonstrating an object manipulation in the scene, followed by robot execution after re-setting the scene.
Task Pi-0 Kinematic DemoDiffusion
Shut Down Laptop 20 10 60
Close Microwave 10 80 90
Drag Basket 10 80 80
Wipe Table 50 0 100
Iron Curtain 20 70 90
Pick up Bear 0 40 60
Pick Place Bowl 0 70 100
Pick Place Banana 0 70 90
Average 13.8 52.5 83.8
@@
TABLE II: Quantitative Results for Real-World Manipulation. Success rates (%) are over 10 trials per task.
Workspace. Fig. 5 describes our real-robot manipulation setup. The scene has 5 cameras including 4 Realsense external cameras and 1 Zed-Mini wrist camera mounted on the robot gripper. We calibrate the four external cameras and use them for triangulating the 3D hand pose so that we can obtain the position and orientation of a human hand in the world coordinates, with origin at the base of the robot. The pre-trained policy Pi-0 requires just two cameras (marked in purple) and the policy does not use any calibration information. While Pi-0 uses Zed 2 camera for external view, we use Realsense camera due to hardware availability, which introduces additional difficulty.
Figure 6: Qualitative comparisons for real-robot manipulation. Rollout progressions (start, intermediate, final frame) for two tasks. Kinematic retargeting (top) produces plausible motion but loses contact before completing the task. Pi-0 (middle) performs general reaching but fails to manipulate the correct object. DemoDiffusion (bottom) reaches the object and maintains contact through task completion. See project page for full videos.
Tasks. We perform evaluations for 8 different manipulation tasks, the language description and the human demonstration for which are described in Fig. 4. Each of these tasks requires a different manipulation strategy involving either prehensile or non-prehensile manipulation. We collect just a single human demonstration per task but for robot execution, we make organic variations in the location of objects in the scene. Additionally, we add other objects in the scene as visual distractors, to see whether DemoDiffusion can generalize under such settings. All variations are consistent across the baselines and our method.
Baselines. For the real-world experiments we compare DemoDiffusion with two baselines. The first is open-loop kinematic retargeting obtained from the human demonstration. The second is directly deploying the Pi-0 model with the language instruction for the respective task. Since Pi-0 operates in the joint velocity space, for DemoDiffusion we first convert the kinematically retargeted trajectory to the joint velocity space with IK and then add noise to the trajectory, followed by denoising with Pi-0. We deploy DemoDiffusion with a heuristically selected noise level $s^{*}/S$ depending on the task – preferring the kinematic trajectory with $s^{*}/S=0.2$ when Pi-0 has a zero success rate and $s^{*}/S=0.4$ otherwise. Note that in the limit of 0 noise added, we recover the first baseline, kinematic retargeting, and in the limit of maximum noise $1$ we obtain the second baseline Pi-0.
Qualitative Visualizations. In Fig. 6, we show rollouts for two tasks across all the methods. We observe that Pi-0 can only perform general reaching behaviors: it either fails to follow the manipulation trajectory needed for the task or confuses the target object with a distractor. On the other hand, we observe that (open-loop) kinematic retargeting from the human demonstration has a plausible motion of the robot end-effector, but it completely misses grasping the iron, and loses contact with the laptop before fully closing it. Compared to these, DemoDiffusion reaches the respective object and also follows the approximate motion of the retargeted trajectory while remaining in contact with the object to solve the tasks.
Shut Laptop Pick Place Banana
DemoDiffusion 60 90
+5cm noise 40 70
with Thumb+Index 30 80
Thumb+Index 0 30
TABLE III: Robustness of DemoDiffusion. We ablate DemoDiffusion under noisy 3D keypoints (2nd row) and with different kinematic retargeting (3rd row). For noisy 3D keypoints, we randomly shift the tracked hand 3D keypoints by 5cm. For different kinematic retargeting, we use the distance between the thumb tip and the index fingertip instead of all fingers.
Quantitative Comparisons. In Table II, we compare DemoDiffusion and the baselines across all the real-world manipulation tasks. DemoDiffusion matches or outperforms both baselines across all 8 tasks, with the largest gains on tasks requiring precise contact – e.g., 60% vs. 10–20% on shutting the laptop and 100% vs. 0–50% on wiping the table. Notably, even for tasks where both baselines fail, DemoDiffusion succeeds by operating at an intermediate noise level $s^{*}$ that combines the trajectory structure from the human demonstration with the policy’s closed-loop refinement.
Noisy Human Demonstrations. To evaluate the robustness of  DemoDiffusion to imperfect inputs, we injected noise into the tracked human trajectories and evaluated DemoDiffusion. Specifically, we place the object at the same location as the human demonstration, and randomly shift the tracked hand 3D keypoints by 5cm. The results in Table III indicate that  DemoDiffusion maintains high-quality performance despite the presence of tracking noise, thereby reducing the reliance on highly accurate 3D annotations in human demonstration.
Different Kinematic Retargeting. Lastly, we further assessed the sensitivity of our method to the choice of retargeting by evaluating an alternative approach (‘Thumb+Index’). Specifically, we use the distance between thumb and index finger [28], instead of all fingers, to infer the grasp. Table III shows the results. While this alternative retargeting itself shows degraded performance compared to the previous retargeting, DemoDiffusion consistently improves task performance despite suboptimal retargeting.
V Discussion
@@
We presented DemoDiffusion, a simple approach to leverage generalist diffusion models for imitating a human demonstration. By retargeting hand motion from the human demonstration into an approximate robot end-effector trajectory and modifying it through de-noising with a pre-trained diffusion policy, DemoDiffusion enables the robot to produce actions that are both aligned with the demonstrated intent and contextually plausible. Our method consistently outperforms both the base diffusion policy and the raw retargeted trajectory across diverse tasks in simulation and the real world, without requiring additional training, paired demonstrations, or reward functions. We envision that this can serve as a starting point for future efforts at human imitation across generic scenarios and perhaps also yield an improved exploration strategy for methods pursuing policy adaptation with online RL.
@@
Our approach has several limitations. First, it assumes that the robot should act similarly to the human to successfully complete the task, which may not hold in scenarios requiring different strategies due to embodiment or environmental constraints. Second, while effective for one-shot imitation (not requiring training / test-time fine-tuning), the method does not produce a reusable task policy that can generalize across arbitrary variations of the task. Third, the quality of the retargeted trajectory is crucial: accurate 3D human motion capture is challenging and errors in retargeting can impact downstream performance, though our method shows a degree of robustness to such noise. Additionally, the method implicitly assumes that the timing and speed of human and robot actions are aligned; extending the approach to allow for temporal alignment at test time is a promising direction of future work.
@@ Acknowledgements
We appreciate the helpful discussions with Yanbo Xu, Qitao Zhao, and Lucas Wu. We would also like to thank Kenny Shaw, Tony Tao, Jiahui Yang, Andrew Wang, Jason Liu, Hengkai Pan, Mohan Kumar Srirama, Peiyuan Yang for helping set up the hardware. This work was supported by gift awards from CISCO, Google, and NSF Award IIS-2442282.
@@ A-A Hardware Setup
A-B Task Setup Description
@@
A-D Inference Details
At each timestep $t$, we index the human demonstration at same timestep $t$, and compute the kinematically retargeted actions. As Pi-0 predicts a chunk of actions with length $H=10$, we also use the action chunk computed from human demonstration from timestep $t$ to t+H1t+H-1. The episode length is set to be the same as the length(number of timesteps) of human demonstrations across all methods.
A-E Kinematic Retargeting Implementation Details
@@ A-F Hyperparameters
TABLE IV: Hyperparameters in Real-world.
@@
A-G Additional Results
Results with Different Hyperparameters. We present the results of DemoDiffusion using two different hyperparameter settings, $s^{*}/S=0.2$ and $s^{*}/S=0.4$. The results are summarized in Table V. Overall, we observe a consistent trend based on the performance of the base diffusion policy (Pi-0):
@@ Appendix B Dexterous Grasping in Simulation
B-A Task Setup Description
@@
TABLE V: Additional Quantitative Results for Real-World Manipulation.
Task Pi-0
Kinematic
Retargeting
DemoDiffusion
($0.2$)
DemoDiffusion
($0.4$)
Shut Laptop 20 10 40 60
Close Microwave 10 80 80 90
Drag Basket 10 80 70 80
Wipe Table 50 0 60 100
Iron Curtain 20 70 70 90
Pick Up Bear 0 40 60 40
Pick Place Banana 0 70 90 20
Pick Place Bowl 0 70 100 0
Average 13.8 52.5 71.3 60.0
@@
B-B Kinematic Retargeting Implementation Details
@@
B-C Robot Policy Implementation Details
@@ B-D Hyperparameters
TABLE VI: Hyperparameters in Simulation.